• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Moffit

    Log entry time 00:40:05 on April 07, 2008

    Entry number 230463

    Followups:

    keyword=Access activity

    Xin and Brad spent some time debugging the trouble with chamber 2. Determined that removal of the 2U13 card allowed for ramping up the full HV (-1600V). Could be a short between a field and signal wire.

    These sets of wires are in the corner of the acceptance.. so not such a big hit.

    Alex reset Right Q2.

    Moved the ribbon cable for 2U'11 to ROC10, Slot 5 ch79-64.